Evaluate the dot product A⃗ ⋅B⃗ if A⃗ =9i^+2j^ and B⃗ =9i^−9j^

Evaluate the dot product A⃗ ⋅B⃗ if A⃗ =9i^−9j^ and B⃗ =9i^+2j^.
Answer #1

here,

A = 9i + 2 j

B = 9 i - 9 j

the dot product , A.B = ( 9 i + 2 j) .(9i - 9 j)

A.B = (81 - 18) units = 63 units

and when

A = 9i - 9 j

B = 9 i + 2 j

the dot product , A.B = ( 9 i - 9 j) .(9i + 2 j)

A.B = (81 - 18) units = 63 units